# Compositional Generative Inverse Design

### 24 Jan 2024 v1

@@ D.2 Surrogate Model for Force Prediction
Figure 7: 54-step trajectories of 2 bodies after performing inverse design using the backpropagation algorithm. Figures (a), (b), (c), and (d) represent the trajectory graphs obtained using GNS, GNS (single step), U-Net, and U-Net (single step) as the forward models, respectively. And (e) is the result of CinDM. The legend of this figure is consistent with Figure 2.
@@ I.2 Influence of initialization
Figure 14: “Re-simulation” error $r_{T}$ of different $T$ in N-body inverse design.
Figure 15: Design performance (lift-to-drag) of different $T$ in 2D airfoil inverse design.
Table 15: Influence of initialization. $r_{T}$ with respect to $T$ for N-body inverse design task. Each number is an of average over 10 batches.
T 2-body 24 steps 2-body 34 steps 2-body 44 steps 2-body 54 steps
@@
Table 16: Influence of initialization. Design performance (lift-tconcerningspect to $T$ for 2D inverse design task. Each number is an of average over 10 batches.
T 1 airfoil 2 airfoils
@@
To analyze the sensitivity of initialization in our approach, we follow a similar methodology discussed in Ren et al. (2020). We consider the “re-simulation” error $r$ of a target objective $y$ as a function of the number of samplings $T$, where each sampling starts from a Gaussian initialization $z$. We use the simulator to obtain the output $\hat{y}$ for each design $x$ from the $T$ design results given the target $y$ and compute the “re-simulation” error $L(\hat{y},y)$. We then calculate the least error among a batch of $T$ design results. This process is repeated for several batches, and the mean least error $r_{T}$ is obtained by averaging over these batches.
@@
Table 15 and Fig 14 present the results for the N-body inverse design task. We consider values of $T$ ranging from 10 to 100, with $N=10$ batches. The target $y$ is set to be 0, which represents the distance to a fixed target point. The results show that $r_{T}$ gradually decreases as $T$ increases in the 24-step design, indicating that the design space is well explored and most solutions can be retrieved even with a small number of samplings. This demonstrates the efficiency of our method in generating designs. Moreover, similar observations can be made when time composition is performed in 34, 44, and 54 steps, indicating the effectiveness of our time composition approach in capturing long-time range physical dependencies and enabling efficient generation in a larger design space.
@@
In the 2D inverse design task, the target $y$ is slightly different. Here, we aim to minimize the model output (drag - lift force). Hence, we adopt the “re-simulation” performance metric, which is the lift/drag ratio, as opposed to the “re-simulation” error used in the N-body task, to evaluate sensitivity to initialization. For each $T$, the lift/drag ratio is chosen as the highest value among the simulation results of a batch of $T$ designed boundaries (or boundary pairs for the 2 airfoils design). Any invalid design results, such as overlapping airfoil pairs in the 2-airfoil design, are removed from the $T$ results before computing the maximal lift/drag ratio. The reported numbers are obtained by averaging over $N$ batches for each $T$.
@@
Table 16 and Fig 15 present the results for the 2D airfoils design task. In the 1 airfoil design column, we observe that the lift/drag performance remains relatively steady for $T\geq 20$. For $T=10$, the lift/drag ratio is relatively low, indicating that the design space is not sufficiently explored due to its high dimensionality (64x64x3 in our boundary mask and offsets representation). In the 2 airfoils design column, the lift/drag ratio increases with $T$. This is attributed to the higher dimensional and more complex design space compared to the single airfoil design task. The stringent constraints on boundary pairs, such as non-overlapping, lead to the presence of complex infeasible regions in the design space. Random initialization may lead to these infeasible regions, resulting in invalid design results. The rate of increase in lift/drag ratio becomes slower when $T\geq 30$, indicating that a majority of solutions have been explored. Despite the training data only containing a single airfoil boundary, which lies in a relatively lower dimensional and simpler design space, our model demonstrates a strong ability to generalize and efficiently generate designs for this challenging 2 body compositional design problem.

### 11 Mar 2024 v2

@@ D.2 Surrogate Model for Force Prediction
Figure 7: 54-step trajectories of 2 bodies after performing inverse design using the CEM algorithm. The trajectory graphs (a), (b), (c), and (d) depict the outcomes using different forward models such as GNS, GNS (single step), U-Net, and U-Net (single step) respectively. Additionally, figure (e) demonstrates the result generated by CinDM. The legend of this figure is consistent with Figure 2.
@@ I.2 Influence of initialization
Figure 14: “Re-simulation” error $r_{B}$ of different $B$ in N-body inverse design.
Figure 15: Design performance (lift-to-drag) of different $B$ in 2D airfoil inverse design.
Table 15: Influence of initialization. $r_{B}$ with respect to $B$ for N-body inverse design task. Each number is an average over 10 batches.
$B$ 2-body 24 steps 2-body 34 steps 2-body 44 steps 2-body 54 steps
@@
Table 16: Influence of initialization. Design performance (lift-to-drag) with respect to $B$ for 2D inverse design task. Each number is an average over 10 batches.
$B$ 1 airfoil 2 airfoils
@@
To analyze the sensitivity of initialization in our approach, we follow a similar methodology discussed in Ren et al. (2020). We consider the “re-simulation” error $r$ of a target objective $y$ as a function of the number of samplings $B$, where each sampling starts from a Gaussian initialization $z$. We use the simulator to obtain the output $\hat{y}$ for each design $x$ from the $B$ design results given the target $y$ and compute the “re-simulation” error $L(\hat{y},y)$. We then calculate the least error among a batch of $B$ design results. This process is repeated for several batches, and the mean least error $r_{B}$ is obtained by averaging over these batches.
@@
Table 15 and Fig 14 present the results for the N-body inverse design task. We consider values of $B$ ranging from 10 to 100, with $N=10$ batches. The target $y$ is set to be 0, which represents the distance to a fixed target point. The results show that $r_{B}$ gradually decreases as $B$ increases in the 24-step design, indicating that the design space is well explored and most solutions can be retrieved even with a small number of samplings. This demonstrates the efficiency of our method in generating designs. Moreover, similar observations can be made when time composition is performed in 34, 44, and 54 steps, indicating the effectiveness of our time composition approach in capturing long-time range physical dependencies and enabling efficient generation in a larger design space.
@@
In the 2D inverse design task, the target $y$ is slightly different. Here, we aim to minimize the model output (drag - lift force). Hence, we adopt the “re-simulation” performance metric, which is the lift/drag ratio, as opposed to the “re-simulation” error used in the N-body task, to evaluate sensitivity to initialization. For each $B$, the lift/drag ratio is chosen as the highest value among the simulation results of a batch of $B$ designed boundaries (or boundary pairs for the 2 airfoils design). Any invalid design results, such as overlapping airfoil pairs in the 2-airfoil design, are removed from the $B$ results before computing the maximal lift/drag ratio. The reported numbers are obtained by averaging over $N=10$ batches for each $B$.
@@
Table 16 and Fig 15 present the results for the 2D airfoils design task. In the 1 airfoil design column, we observe that the lift/drag ratio is relatively low for $B=10$, indicating that the design space is not sufficiently explored due to its high dimensionality ($64\times 64\times 3$ in our boundary mask and offsets representation). For $B\geq 20$, the lift/drag performance remains steady. In the 2 airfoils design column, the lift/drag ratio increases roughly with $B$. This is attributed to the higher dimensional and more complex design space compared to the single airfoil design task. The stringent constraints on boundary pairs, such as non-overlapping, lead to the presence of complex infeasible regions in the design space. Random initialization may lead to these infeasible regions, resulting in invalid design results. The rate of increase in lift/drag ratio becomes slower when $B\geq 30$, indicating that a majority of solutions have been explored. Despite the training data only containing a single airfoil boundary, which lies in a relatively lower dimensional and simpler design space, our model demonstrates a strong ability to generalize and efficiently generate designs for this challenging 2 body compositional design problem.